Vitex, Inc. manufactures a popular consumer product and it has provided the following data excerpts from its standard cost system:

Inputs (1) Standard Quantity or Hours (2)
Standard
Price
or Rate
Standard
Cost
(1) × (2)
Direct materials 2.40 pounds $ 16.30 per pound $ 39.12
Direct labor 1.00 hours $ 15.90 per hour $ 15.90
Variable manufacturing overhead 1.00 hours $ 9.60 per hour $ 9.60
Total standard cost per unit $ 64.62
Total Variances Reported
Standard
Cost*
Price
or Rate
Quantity or
Efficiency
Direct materials $ 625,920 $ 11,716 F $ 32,600 U
Direct labor $ 254,400 $ 3,400 U $ 15,900 U
Variable manufacturing overhead $ 153,600 $ 4,800 F $ ? U

*Applied to Work in Process during the period.

The company's manufacturing overhead cost is applied to production on the basis of direct labor-hours. All of the materials purchased during the period were used in production. Work in process inventories are insignificant and can be ignored.

Required:

1. How many units were produced last period?

2. How many pounds of direct material were purchased and used in production?

3. What was the actual cost per pound of material? (Round your answer to 2 decimal places.)

4. How many actual direct labor-hours were worked during the period?

5. What was the actual rate paid per direct labor-hour? (Round your answer to 2 decimal places.)

6. How much actual variable manufacturing overhead cost was incurred during the period?
Answer #1

Answer 1.

Direct Materials:

Standard costs = Standard cost per unit * Actual units produced
$625,920 = $39.12 * Actual units produced
Actual units produced = 16,000

Answer 2.

Standard materials allowed = Standard materials per unit * Actual units produced
Standard materials allowed = 2.40 * 16,000
Standard materials allowed = 38,400 pounds

Material quantity variance = Standard materials price per pound * (Actual materials used - Standard materials allowed)
$32,600 = $16.30 * (Actual materials used - 38,400)
2,000 = Actual materials used - 38,400
Actual materials used = 40,400 pounds

Answer 3.

Material price variance = Actual materials used * (Actual materials price per pound - Standard materials price per pound)
-$11,716 = 40,400 * (Actual materials price per pound - $16.30)
-$0.29 = Actual materials price per pound - $16.30
Actual materials price per pound = $16.01

Answer 4.

Standard labor hours allowed = Standard labor hours per unit * Actual units produced
Standard labor hours allowed = 1.00 * 16,000
Standard labor hours allowed = 16,000

Labor efficiency variance = Standard labor rate per hour * (Actual labor hours worked - Standard labor hours allowed)
$15,900 = $15.90 * (Actual labor hours worked - 16,000)
1,000 = Actual labor hours worked - 16,000
Actual labor hours worked = 17,000

Answer 5.

Labor rate variance = Actual labor hours worked * (Actual labor rate per hour - Standard labor rate per hour)
$3,400 = 17,000 * (Actual labor rate per hour - $15.90)
$0.20 = Actual labor rate per hour - $15.90
Actual labor rate per hour = $16.10

Answer 6.

Variable manufacturing overhead rate variance = Actual variable manufacturing overhead cost - Actual labor hours worked * Standard variable manufacturing overhead rate per hour
-$4,800 = Actual variable manufacturing overhead cost - 17,000 * $9.60
-$4,800 = Actual variable manufacturing overhead cost - $163,200
Actual variable manufacturing overhead cost = $158,400